A television set that originally cost $1955 was reduced in price by successive discounts of 5%, 8% and 24%. What was the equivalent single rate of discount?

Answer by ikleyn(52775)

.
A television set that originally cost $1955 was reduced in price by successive discounts
of 5%, 8% and 24%. What was the equivalent single rate of discount?
~~~~~~~~~~~~~~~~~~~~~

Let X be the original cost, "e" be the equivalent discount for these three 
successive discounts, as a decimal value.


Then 

    (1-e)X = (1-0.05)*(1-0.08)*(1-0.24)X.


Cancel X in both sides.  You will get

    1 - e = 0.95*0.92*0.76 = 0.66424,

    e = 1 - 0.66424 = 0.33576.


The equivalent discount is 0.33576,  or  33.576%.    ANSWER

Solved.

The answer does not depend on the original cost value, which is $1995 in this problem.